# FedFACT: A Provable Framework for Controllable Group-Fairness Calibration in Federated Learning

### v1

@@ B.6 Proof of Theorem 5
[5] The subdifferential of the function $F(x)=\mathbb{E}\{f(x,\omega)\}$ at a point $x$ is given by
@@
[…]
@@
[59] Let $f_{1},\dots,f_{m}:$ $\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be convex functions. Define $f(x)=\max\{f_{1}(x),\dots,f_{m}(x)\},\quad\forall x\in\mathbb{R}^{n}$. For $x_{0}\in\bigcap_{i=1}^{m}\text{dom}f_{i}$, define $I(x_{0})=\{i\mid f_{i}(x_{0})=f(x_{0})\}$. Then $\partial f(x_{0})=\text{conv}\bigcup_{i\in I(x_{0})}\partial f_{i}(x_{0}).$
@@
[59] Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a convex continuous function. We consider the minimizer $x^{*}$ of the function $f$ over the set $B$. Then for $x^{*}$ to be locally optimal it is necessary that
@@
where $\mathbf{M}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{p_{a,k}}\Big{[}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}-\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\mathbf{D}^{a,k}_{u_{l,k}}\Big{]}$. Considering the personalized attribute-aware classifier $h_{k}(x,a),k\in[N]$ in post-processing, the inner function turns to
@@
$\begin{split}\min_{\lambda,\mu}H(\lambda,\mu):=&\sum_{k\in[N]}\sum_{a\in\mathcal{A}}p_{a,k}\underset{X\sim\mathcal{P}^{X}_{a,k}}{\mathbb{E}}\Big{[}\max_{y\in[m]}\Big{(}\big{[}\mathbf{M}^{\lambda,\mu}(a,k)\big{]}^{\top}\eta(X,a,k)\Big{)}_{y}\Big{]}+\xi^{g}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})\\ &\qquad+\sum_{k\in[N]}\xi^{l,k}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}+\mu^{(2)}_{k,u_{l,k}}).\end{split}$ (42)
@@
where $\mathbf{h}^{fair}$ denotes a classifier that satisfies fairness constraints for given $\xi^{g},\xi^{l,k}>0$. Hence, we arrive at
@@
holds for all $\lambda,\mu\geq 0$. Notice that given $\lambda,\mu\geq 0$, this inequality holds for any $\xi^{g},\xi^{l,k}>0$. Let $\xi^{g}\to 0,\xi^{l,k}\to 0$, combining (42) and (43) gives that
@@
Therefore, the dual problem has lower bound
$H(\lambda,\mu)\geq\xi^{g}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})+\sum_{k\in[N]}\xi^{l,k}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}+\mu^{(2)}_{k,u_{l,k}})$ (45)
@@
With a slight abusion of notation, let score function $f(x,a,k)=\big{[}\mathbf{M}^{\lambda,\mu}(a,k)\big{]}^{\top}\eta(x,a,k)$, by Lemma B.5, we have
@@
[…]
@@
where $B^{t}_{y}:=\{\exists t\neq y,f_{t}(x,a,k)\geq f_{i}(x,a,k),\forall i\in[m];f_{t}(x,a,k)=f_{y}(x,a,k)\}$ with $b_{t}\in[0,1]$. Since the convex hull is a interval here, by Caratheodory’s theorem, it can be characterized by two point here (the initial point $e_{y}$ and another point $e_{t}$ in the convex hull). Without loss of generality, we assume the existence of one $e_{t}$ such that $f_{t}(x,a,k)=f_{y}(x,a,k)$ here. We know that $f_{t}(x,a,k)-f_{y}(x,a,k)=[\eta(x,a,k)]^{\top}\mathbf{M}^{\lambda,\mu}(a,k)(e^{t}-e^{y})$. With Asumption 1, we obtain that the measure of $B^{t}_{y}$ is 00, unless the $t$-th and $y$-th column of $\mathbf{M}^{\lambda,\mu}(a,k)$ are equal. An effective simplification is to exclude all $\lambda^{\prime},\mu^{\prime}$ that cause $\mathbf{M}^{\lambda^{\prime},\mu^{\prime}}(a,k)(e^{t}-e^{y})=0$. Since we suppose that the non-zero columns of each $\mathbf{D}^{a,k}_{u}$ are distinct, the dual parameter $\lambda^{\prime},\mu^{\prime}\in S_{t,y}$, such that $\mathbf{M}^{\lambda^{\prime},\mu^{\prime}}(a,k)(e^{t}-e^{y})=0$, constructs the empty relative interior in the dual parameter space. By the convexity of the objective function, we have $\inf_{\lambda,\mu\notin S_{t,y}}H(\lambda,\mu)=\min_{\lambda,\mu}H(\lambda,\mu)$, due to the density of $(\lambda,\mu)\notin S_{t,y}$.
@@
[…]
@@
[…]
@@
The local fairness guarantee also can be derived from the optimality of $\mu^{*}$. The proof techniques are extremely similar to our proof with respect to $\lambda^{*}$. Hence, we omit the proof of the local fairness guarantee here. The result turns out that $|\mathscr{D}^{l,k}_{u_{l},k}(\mathbf{h}^{\lambda,\mu^{*}})|\leq\xi^{l,k},k\in[N]$.
@@
which satisfies the optimality conditions for the dual solution of the constrained optimization problem. The same holds for the local fairness constraints $\mathscr{D}^{l,k}(\mathbf{h^{\lambda^{*},\mu^{*}}})$. Consequently, the Lagrangian function equals to risk function when plugging in optimal classifier, $\mathcal{L}(h^{\lambda^{*},\mu^{*}},\lambda^{*},\mu^{*})=\mathcal{R}(h^{\lambda^{*},\mu^{*}})$. For any other classifiers $\mathbf{h}^{\prime}$ that satisfies the global and local fairness constraints, denoting its corresponding dual parameter to maximize the outer problem as $\lambda^{\prime},\mu^{\prime}$, it can be deduced that
@@ B.7 Proof of Prosition 6
Note that $\lambda\in\mathbb{R}^{2|\mathcal{U}_{g}|}_{\geq 0}$ and $\mu_{k}\in\mathbb{R}^{2|\mathcal{U}_{l,k}|}_{\geq 0}$, the operator $\|\cdot\|_{1}$ is linear in dual parameters’ domain. We can just write
@@
$+\frac{\xi^{l,k}}{\hat{p}_{k}}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu_{k,u_{l,k}}^{(1)}+\mu_{k,u_{l,k}}^{(2)}),$
where $\widehat{\mathbf{M}}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{\hat{p}_{a,k}}\Big{[}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\widehat{\mathbf{D}}^{a,k}_{u_{g}}-\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\widehat{\mathbf{D}}^{a,k}_{u_{l,k}}\Big{]}$ and $\sigma_{\beta}(x)=\sum_{i=1}^{m}\frac{\exp(x_{i}/\beta)}{\sum_{j=1}^{m}\exp(x_{j}/\beta)}x_{i}$.
@@
Set $f_{i}(\lambda)\;:=\;\sigma_{\beta}\bigl{(}z_{i}(\lambda)\bigr{)}.$ and let $f_{i}(\lambda)=\sigma_{\beta}\bigl{(}z_{i}(\lambda)\bigr{)},\qquad\sigma_{\beta}(x)=\sum_{j=1}^{m}\frac{e^{x_{j}/\beta}}{\sum_{\ell=1}^{m}e^{x_{\ell}/\beta}}\,x_{j},$ By the chain rule and second-order derivatives, $\nabla_{\lambda}f_{i}(\lambda)=A_{i}^{\top}\nabla_{x}\sigma_{\beta}\bigl{(}z_{i}(\lambda)\bigr{)},\qquad\nabla^{2}_{\lambda}f_{i}(\lambda)=A_{i}^{\top}\bigl{[}\nabla^{2}\sigma_{\beta}\bigl{(}z_{i}(\lambda)\bigr{)}\bigr{]}A_{i}.$ Hence, due to the boundedness of $\|\lambda\|_{1}$, the inside $z_{i}(\lambda)$ is bounded, setting the upper bound as $R$ for simplification here, $\bigl{\|}\nabla^{2}_{\lambda}f_{i}(\lambda)\bigr{\|}_{2}\;\leq\;\|A_{i}\|_{2}^{2}\;\sup_{x}\bigl{\|}\nabla^{2}\sigma_{\beta}(x)\bigr{\|}_{2}=\|A_{i}\|_{2}^{2}\,L_{\sigma},$ showing $f_{i}$ is $\|A_{i}\|_{2}^{2}L_{\sigma}$-smooth. The linear term in $\lambda$ has zero Hessian. Therefore, since the average of smooth functions is smooth with averaged constants, the function $\widehat{H}_{k}^{\prime}(\lambda,\mu_{k})$ is $L$-smooth in $\lambda$ with $L=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\|A_{i}\|_{2}^{2}\,L_{\sigma}.$ Following the similar proof procedure, we can obtain the smoothness of $\widehat{H}_{k}^{\prime}(\lambda,\mu_{k})$ to $\mu_{k}$. $\square$
@@ B.8 Generalization Error For Post-processing Algorithm
If classifiers $\widehat{\mathbf{h}}^{*}=(\widehat{h}^{*}_{1},\dots,\widehat{h}^{*}_{k})$ with dual parameters $(\widehat{\lambda}^{*},\widehat{\mu}^{*})$ form an optimal solution of the empirical plug-in estimation of (5), denoting $\rho(n,\delta)=\sqrt{\frac{8|\mathcal{A}|m^{2}\log(n+1)}{n}}+\sqrt{\frac{2\log(m^{2}|\mathcal{A}|N/\delta)}{n}}$, $B_{g}=\max_{a\in\mathcal{A},k\in[N]}\|{\mathbf{D}}^{a,k}_{u_{g}}\|_{1},\widehat{B}_{g}=\max_{a\in\mathcal{A},k\in[N]}\|\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{1},\Omega_{n}^{g}=\max_{a\in\mathcal{A},k\in[N]}\|\mathbf{D}^{a,k}_{u_{g}}-\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{\infty}$, and $B_{l,k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{l,k}}\|_{1},\widehat{B}_{l,k}=\max_{a\in\mathcal{A}}\|\widehat{\mathbf{D}}^{a,k}_{u_{l,k}}\|_{1},\Omega_{n}^{l,k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{l,k}}-\widehat{\mathbf{D}}^{a,k}_{u_{l,k}}\|_{\infty},k\in[N]$.
(1) Let $10<\delta<$ 1, suppose that $n>\frac{2|\mathcal{A}|N\widehat{B}_{l,k}}{p_{\min}\xi^{g}}+\frac{1}{2p_{\min}^{2}}\log\frac{1}{\delta}$, then with probability at least $1-2|\mathcal{A}|\delta$,
@@
(2) Let $10<\delta<$ 1, suppose that $n>\frac{2|\mathcal{A}|\widehat{B}_{g}}{p_{\min}\xi^{l,k}}+\frac{1}{2p_{\min}^{2}}\log\frac{1}{\delta}$, then with probability at least $1-2|\mathcal{A}|\delta$,
$|{\mathscr{D}}^{l,k}({\widehat{\mathbf{h}}^{*}})|$ $\leq\xi^{l,k}+\mathcal{O}(|\mathcal{A}|B_{l,k}\rho(n,N\delta/|\mathcal{A}||\mathcal{U}_{g}|))+\Omega^{l,k}_{n}+\frac{|\mathcal{A}|\widehat{B}_{l,k}}{n_{\min}},\ k\in[N].$
@@
[…]
Proof of Lemma B.7. By Hoeffding’s inequality, for any $t>0$, $\mathbb{P}\bigl{(}S_{n}-\mathbb{E}[S_{n}]\leq-t\bigr{)}\;\leq\;\exp\!\Bigl{(}-\frac{2\,t^{2}}{n}\Bigr{)}.$ Since $\mathbb{E}[S_{n}]=np$, set $t=np-M.$ Then

### v2

@@ B.6 Proof of Theorem 5
[lemma-subgradient-Bertsekas1973StochasticOP] The subdifferential of the function $F(x)=\mathbb{E}\{f(x,\omega)\}$ at a point $x$ is given by
@@
$\begin{split}\mathbb{E}\{\partial f(x,&\omega)\}:=\int_{\Omega}\partial f(x,\omega)d\mathbb{P}(\omega)\\ &=\bigg\{x^{*}\in\mathbb{R}^{n}\mid x^{*}=\int_{\Omega}x^{*}(\omega)d\mathbb{P}(\omega),x^{*}(\cdot):\text{measurable},x^{*}(\omega)\in\partial f(x,\omega)\text{ a.e.}\bigg\}.\end{split}$
@@
[Variational-Analysis-RockWets98] Let $f_{1},\dots,f_{m}:$ $\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be convex functions. Define $f(x)=\max\{f_{1}(x),\dots,f_{m}(x)\},\quad\forall x\in\mathbb{R}^{n}$. For $x_{0}\in\bigcap_{i=1}^{m}\text{dom}f_{i}$, define $I(x_{0})=\{i\mid f_{i}(x_{0})=f(x_{0})\}$. Then $\partial f(x_{0})=\text{conv}\bigcup_{i\in I(x_{0})}\partial f_{i}(x_{0}).$
@@
[Variational-Analysis-RockWets98] Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a convex continuous function. We consider the minimizer $x^{*}$ of the function $f$ over the set $B$. Then for $x^{*}$ to be locally optimal it is necessary that
@@
where $\mathbf{M}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{p_{a,k}}\Big[\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}-\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\mathbf{D}^{a,k}_{u_{k}}\Big]$. Considering the personalized attribute-aware classifier $h_{k}(x,a),k\in[N]$ in post-processing, the inner function turns to
@@
Before exploring the optimal solution of outer optimization, we first prove that the optimal dual parameter $\lambda^{*}\in\mathbb{R}^{2|\mathcal{U}_{g}|}_{\geq 0},\mu^{*}\in\mathbb{R}^{2\sum_{k=1}^{N}|\mathcal{U}_{k}|}_{\geq 0}$ is bounded. Define the Hilbert space on $\mathcal{F}:=\{f:\mathcal{X}\to\mathbb{R}^{m}\}$ with inner product $\langle f,g\rangle=\int_{\mathcal{X}}f^{\top}gd\mathcal{P}(x)$. Then the classifier space $\mathcal{H}:{\mathcal{X}\to\Delta_{m}}$ is a convex subset of $\mathcal{F}$. Therefore, we can also consider the topology structure on $\mathcal{H}$ or $\mathcal{H}^{|\mathcal{A}|}$. Since we assume that $\forall\xi^{g},\xi^{k}>0$, the feasible set of the primal problem is non-empty, it indicates that the feasible set of the primal problem has non-empty interior for any positive $\xi^{g},\xi^{k}$. It is clear that for $\forall\xi^{g},\xi^{k}>0$, the dual problem
@@
where $\mathbf{h}^{fair}$ denotes a classifier that satisfies fairness constraints for given $\xi^{g},\xi^{k}>0$. Hence, we arrive at
@@
holds for all $\lambda,\mu\geq 0$. Notice that given $\lambda,\mu\geq 0$, this inequality holds for any $\xi^{g},\xi^{k}>0$. Let $\xi^{g}\to 0,\xi^{k}\to 0$, combining (42) and (43) gives that
@@
Therefore, the dual problem has a lower bound
$H(\lambda,\mu)\geq\xi^{g}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})+\sum_{k\in[N]}\xi^{k}\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}+\mu^{(2)}_{k,u_{k}})$ (45)
@@
With a slight abuse of notation, let score function $f(x,a,k)=\big[\mathbf{M}^{\lambda,\mu}(a,k)\big]^{\top}\eta(x,a,k)$, by Lemma B.5, we have
@@
$=\frac{1}{p_{a,k}}\sum_{y\in[m]}\int_{\{x:h_{k}^{\lambda,\mu}(x,a)=y\}}\left[\ \mathrm{conv}\left(\bigcup_{i\in\arg\max_{i}(f_{i}(x,a,k))}-[\eta(x,a,k)]^{T}\mathbf{D}^{a,k}_{u_{g}}e_{y}\right)\right]d\mathcal{P}^{X}_{a,k}(x)$ (49)
@@
where $B^{t}_{y}:=\{\exists t\neq y,f_{t}(x,a,k)\geq f_{i}(x,a,k),\forall i\in[m];f_{t}(x,a,k)=f_{y}(x,a,k)\}$ with $b_{t}\in[0,1]$. Since the convex hull is a interval here, by Caratheodory’s theorem, it can be characterized by two point here (the initial point $e_{y}$ and another point $e_{t}$ in the convex hull). Without loss of generality, we assume the existence of one $e_{t}$ such that $f_{t}(x,a,k)=f_{y}(x,a,k)$ here. We know that $f_{t}(x,a,k)-f_{y}(x,a,k)=[\eta(x,a,k)]^{\top}\mathbf{M}^{\lambda,\mu}(a,k)(e^{t}-e^{y})$. With Asumption 1, we obtain that the measure of $B^{t}_{y}$ is 0, unless the $t$-th and $y$-th column of $\mathbf{M}^{\lambda,\mu}(a,k)$ are equal. An effective simplification is to exclude all $\lambda^{\prime},\mu^{\prime}$ that cause $\mathbf{M}^{\lambda^{\prime},\mu^{\prime}}(a,k)(e^{t}-e^{y})=0$. Since we suppose that the non-zero columns of each $\mathbf{D}^{a,k}_{u}$ are distinct, the dual parameter $\lambda^{\prime},\mu^{\prime}\in S_{t,y}$, such that $\mathbf{M}^{\lambda^{\prime},\mu^{\prime}}(a,k)(e^{t}-e^{y})=0$, constructs the empty relative interior in the dual parameter space. By the convexity of the objective function, we have $\inf_{\lambda,\mu\notin S_{t,y}}H(\lambda,\mu)=\min_{\lambda,\mu}H(\lambda,\mu)$, due to the density of $(\lambda,\mu)\notin S_{t,y}$.
@@
$\frac{\partial}{\partial\lambda^{(1)}_{u_{g}}}H(\lambda,\mu)$ $=\sum_{k\in[N]}\sum_{a\in\mathcal{A}}\sum_{y\in[m]}\int_{\{x:h_{k}^{\lambda,\mu}(x,a)=y\}}\Big[-\Big(\big[\mathbf{D}^{a,k}_{u_{g}}\big]^{\top}\eta(x,a,k)\Big)_{y}\Big]d\mathcal{P}^{X}_{a,k}(x)+\xi^{g}$
@@
$\frac{\partial}{\partial\lambda^{(2)}_{u_{g}}}H(\lambda,\mu)$ $=\sum_{k\in[N]}\sum_{a\in\mathcal{A}}\int_{\mathcal{X}}\big[\eta(x,a,k)\big]^{\top}\mathbf{D}^{a,k}_{u_{g}}h^{\lambda,\mu}_{k}(x,a)d\mathcal{P}^{X}_{a,k}(x)+\xi^{g}$
@@
The local fairness guarantee also can be derived from the optimality of $\mu^{*}$. The proof techniques are extremely similar to our proof with respect to $\lambda^{*}$. Hence, we omit the proof of the local fairness guarantee here. The result turns out that $|\mathscr{D}^{k}_{u_{k}}(\mathbf{h}^{\lambda,\mu^{*}})|\leq\xi^{k},k\in[N]$.
@@
which satisfies the optimality conditions for the dual solution of the constrained optimization problem. The same holds for the local fairness constraints $\mathscr{D}^{k}(\mathbf{h^{\lambda^{*},\mu^{*}}})$. Consequently, the Lagrangian function equals to risk function when plugging in optimal classifier, $\mathcal{L}(h^{\lambda^{*},\mu^{*}},\lambda^{*},\mu^{*})=\mathcal{R}(h^{\lambda^{*},\mu^{*}})$. For any other classifiers $\mathbf{h}^{\prime}$ that satisfies the global and local fairness constraints, denoting its corresponding dual parameter to maximize the outer problem as $\lambda^{\prime},\mu^{\prime}$, it can be deduced that
@@ B.7 Proof of Prosition 6
Note that $\lambda\in\mathbb{R}^{2|\mathcal{U}_{g}|}_{\geq 0}$ and $\mu_{k}\in\mathbb{R}^{2|\mathcal{U}_{k}|}_{\geq 0}$, the operator $\|\cdot\|_{1}$ is linear in dual parameters’ domain. We can just write
@@
$+\frac{\xi^{k}}{\hat{p}_{k}}\sum_{u_{k}\in\mathcal{U}_{k}}(\mu_{k,u_{k}}^{(1)}+\mu_{k,u_{k}}^{(2)}),$
where $\widehat{\mathbf{M}}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{\hat{p}_{a,k}}\Big[\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\widehat{\mathbf{D}}^{a,k}_{u_{g}}-\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\widehat{\mathbf{D}}^{a,k}_{u_{k}}\Big]$ and $\sigma_{\beta}(x)=\sum_{i=1}^{m}\frac{\exp(x_{i}/\beta)}{\sum_{j=1}^{m}\exp(x_{j}/\beta)}x_{i}$.
@@
Set $f_{i}(\lambda)\;:=\;\sigma_{\beta}\bigl(z_{i}(\lambda)\bigr).$ and let $f_{i}(\lambda)=\sigma_{\beta}\bigl(z_{i}(\lambda)\bigr),\qquad\sigma_{\beta}(x)=\sum_{j=1}^{m}\frac{e^{x_{j}/\beta}}{\sum_{\ell=1}^{m}e^{x_{\ell}/\beta}}\,x_{j},$ By the chain rule and second‐order derivatives, $\nabla_{\lambda}f_{i}(\lambda)=A_{i}^{\top}\nabla_{x}\sigma_{\beta}\bigl(z_{i}(\lambda)\bigr),\qquad\nabla^{2}_{\lambda}f_{i}(\lambda)=A_{i}^{\top}\bigl[\nabla^{2}\sigma_{\beta}\bigl(z_{i}(\lambda)\bigr)\bigr]A_{i}.$ Hence, due to the boundedness of $\|\lambda\|_{1}$, the inside $z_{i}(\lambda)$ is bounded, setting the upper bound as $R$ for simplification here, $\bigl\|\nabla^{2}_{\lambda}f_{i}(\lambda)\bigr\|_{2}\;\leq\;\|A_{i}\|_{2}^{2}\;\sup_{x}\bigl\|\nabla^{2}\sigma_{\beta}(x)\bigr\|_{2}=\|A_{i}\|_{2}^{2}\,L_{\sigma},$ showing $f_{i}$ is $\|A_{i}\|_{2}^{2}L_{\sigma}$‐smooth. The linear term in $\lambda$ has zero Hessian. Therefore, since the average of smooth functions is smooth with averaged constants, the function $\widehat{H}_{k}^{\prime}(\lambda,\mu_{k})$ is LL‐smooth in $\lambda$ with $L=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\|A_{i}\|_{2}^{2}\,L_{\sigma}.$ Following the similar proof procedure, we can obtain the smoothness of $\widehat{H}_{k}^{\prime}(\lambda,\mu_{k})$ to $\mu_{k}$. $\square$
@@ B.8 Generalization Error For Post-Processing Algorithm
If classifiers $\widehat{\mathbf{h}}^{*}=(\widehat{h}^{*}_{1},\dots,\widehat{h}^{*}_{k})$ with dual parameters $(\widehat{\lambda}^{*},\widehat{\mu}^{*})$ form an optimal solution of the empirical plug-in estimation of (7), denoting $\rho(n,\delta)=\sqrt{\frac{8|\mathcal{A}|m^{2}\log(n+1)}{n}}+\sqrt{\frac{2\log(m^{2}|\mathcal{A}|N/\delta)}{n}}$, $B_{g}=\max_{a\in\mathcal{A},k\in[N]}\|{\mathbf{D}}^{a,k}_{u_{g}}\|_{1},\widehat{B}_{g}=\max_{a\in\mathcal{A},k\in[N]}\|\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{1},\Omega_{n}^{g}=\max_{a\in\mathcal{A},k\in[N]}\|\mathbf{D}^{a,k}_{u_{g}}-\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{\infty}$, and $B_{k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{k}}\|_{1},\widehat{B}_{k}=\max_{a\in\mathcal{A}}\|\widehat{\mathbf{D}}^{a,k}_{u_{k}}\|_{1},\Omega_{n}^{k}=\max_{a\in\mathcal{A}}\|\mathbf{D}^{a,k}_{u_{k}}-\widehat{\mathbf{D}}^{a,k}_{u_{k}}\|_{\infty},k\in[N]$.
(1) Let $0<\delta<1$, suppose that $n>\frac{2|\mathcal{A}|N\widehat{B}_{k}}{p_{\min}\xi^{g}}+\frac{1}{2p_{\min}^{2}}\log\frac{1}{\delta}$, then with probability at least $1-2|\mathcal{A}|\delta$,
@@
(2) Let $0<\delta<1$, suppose that $n>\frac{2|\mathcal{A}|\widehat{B}_{g}}{p_{\min}\xi^{k}}+\frac{1}{2p_{\min}^{2}}\log\frac{1}{\delta}$, then with probability at least $1-2|\mathcal{A}|\delta$,
$|{\mathscr{D}}^{k}({\widehat{\mathbf{h}}^{*}})|$ $\leq\xi^{k}+\mathcal{O}(|\mathcal{A}|B_{k}\rho(n,N\delta/|\mathcal{A}||\mathcal{U}_{g}|))+\Omega^{k}_{n}+\frac{|\mathcal{A}|\widehat{B}_{k}}{n_{\min}},\ k\in[N].$
@@
Let $X_{1},\dots,X_{n}$ be independent $\mathrm{Bernoulli}(p)$ random variables and define $S_{n}\;=\;\sum_{i=1}^{n}X_{i}.$ Fix any $M\in(0,np)$ and confidence level $\delta\in(0,1)$. If the sample size satisfies $n\;\geq\;\frac{2M}{p}\;+\;\frac{1}{2p^{2}}\,\log\frac{1}{\delta},$ then we have $\mathbb{P}\bigl(S_{n}>M\bigr)\;\geq\;1-\delta.$
Proof of Lemma B.7. By Hoeffding’s inequality, for any $t>0$, $\mathbb{P}\bigl(S_{n}-\mathbb{E}[S_{n}]\leq-t\bigr)\;\leq\;\exp\!\Bigl(-\frac{2\,t^{2}}{n}\Bigr).$ Since $\mathbb{E}[S_{n}]=np$, set t=npM.t=np-M. Then